OpenStudy (dranz10):

The population f(x), in millions, of State A of a country after x years is represented by the function shown below: f(x) = (1.08)x The graph shows the population g(x), in millions, of State B of the country after x years:

OpenStudy (dranz10):

Its actually, f(x)= (1.08)^x

OpenStudy (dranz10):

The choices are... Which conclusion is correct about the population of State A and State B? The original population of State B was half of the original population of State A. The original population of State A was half of the original population of State B. The original population of State B was one-fourth of the original population of State A. The original population of State A was one-fourth of the original population of State B.

OpenStudy (dranz10):

@agent0smith @jhonyy9

OpenStudy (mathmale):

Not your fault, but I find the graph to be incorrect. If x=0 in f(x)= (1.08)^x, we get f(0)=(1.08)^0=1 (not 2, as the graph indicates). Might be better to move on to the next problem.